Each of 100 firms in a perfectly competitive industry has an AVC function AVC= 2Q where Q represents the output produced. If the price of the product is Rs.100 find the total quantity supplied in the market when a) total fixed cost is 10,000 and b) total fixed cost is Rs.20,000.


1
Expert's answer
2021-02-22T15:12:08-0500

If the average variable cost for each firm is AVC=2q, then the total variable cost is:

"TVC=2q^2"

And the marginal cost is:

"MC=\dfrac{d TVC}{d q}=4q"

For a competitive firm, P=MC and the marginal cost is the supply curve for an individual firm. Thus, the supply curve for a single firm is:

"P=4q\\[0.3cm]
q=0.25P"

There are 100 firms in the industry. Therefore, the industry supply curve is:

"Q=nq\\[0.3cm]
Q=100(0.25P)\\[0.3cm]
Q=25P"

When the market price is P=100, the total quantity supplied is:

"Q=25(100)\\[0.3cm]
Q=\boxed{2500}"

The change in the fixed cost will not affect the market supply since the fixed costs are not used in decision making.